Integration by Parts Fractions

∫dy/(y^2+2y)


1
Expert's answer
2021-04-27T08:51:47-0400

Let

"\frac{1}{{{y^2} + 2y}} = \frac{1}{{y(y + 2)}} = \frac{A}{y} + \frac{B}{{y + 2}} = \frac{{Ay + 2A + By}}{{y(y + 2)}} = \frac{{\left( {A + B} \right)y + 2A}}{{y(y + 2)}}"

Then

"\left\{ \begin{array}{l}
A + B = 0\\
2A = 1
\end{array} \right."

"\left\{ \begin{array}{l}
A = \frac{1}{2}\\
B = - \frac{1}{2}
\end{array} \right."

Then

"\int {\frac{{dy}}{{{y^2} + 2y}} = \frac{1}{2}} \int {\frac{{dy}}{y}} - \frac{1}{2}\int {\frac{{dy}}{{y + 2}} = \frac{1}{2}} \ln |y| - \frac{1}{2}\ln |y + 2| + C"